# DESI Data Release 1: Stellar Catalogue

### v1

@@ 4.1 Distribution of stellar parameters and abundances
Figure 5 shows the distribution of surface gravity vs (Kiel diagram) using measurements from both stellar pipelines. We show the stars observed in the main survey with signal-to-noise in the R arm of the instrument above 10. The main improvement in this distribution compared to the EDR catalogue (c.f. Figure 9 in K24) is the disappearance of clustering near the grid-points in and for the RVS catalogue. Otherwise, the distribution shows the expected main sequence, red giant branch as well as horizontal branch. For dwarf stars () with lower effective temperatures ( K), the surface gravity is not a smooth function of temperature, similar to what was seen in EDR. The morphology of the diagrams for RVS and SP pipeline is similar, but the SP parameters show some gridding of temperatures for cool stars, also the RVS red giant branch seem less populated at .
@@
Finally, Figure 7 shows the metallicity distribution for stars in the main survey. We use the same selection of stars as for the previous figure. The distributions clearly show the peaks associated with the MW disk, and Gaia Sausage-Enceladus at [Fe/H] (see, e.g., Belokurov et al., 2018; Helmi et al., 2018; Feuillet et al., 2020). We note that although the metallicity distribution shows a almost power-law-like behaviour from [Fe/H] to lower metallicities, it also shows small "wiggles", for example at [Fe/H], where the metallicity distribution (as measured by two pipelines) briefly flattens. That could be the result of debris deposition from a relatively massive progenitor with the mean metallicity of (i.e. something like Thamnos or I’itoi; see Naidu et al., 2020). Overall the distribution of RVS and SP metallicities show similar behaviour, except near the edges. At high metallicities the RVS sample have more stars than SP (this is mainly caused by the exclusion of SP measurements fitted by PHOENIX models). At low metallicities we can see some differences in the SP and RVS distributions of [Fe/H] mostly due to the fact that the PHOENIX grid used by the RVS pipeline has a limit at [Fe/H], while the SP pipeline has a limit at [Fe/H]. We note that the SP metallicities show an excess near the grid edge at [Fe/H]= which is not physical. Using the same sample as shown in the figure, we can estimate the total number of extremely metal-poor stars (i.e with [Fe/H]) in the bright program of the main survey – it is if we use RVS abundances, and 3000 with SP abundances. These numbers rise by 2,000 if we include all surveys and programs.
@@ 4.2 Comparison of abundances to external measurements
In the previous section we looked at the overall distribution of stellar parameters and abundances. In this section we focus on iron abundance and compare our results to those from other large medium- and high-resolution surveys such as APOGEE (Majewski et al., 2017), GALAH (De Silva et al., 2015) and Gaia RVS (Cropper et al., 2018). We use the data from the DR17 catalogue (Abdurro’uf et al., 2022) of APOGEE, the DR4 catalogue (Buder et al., 2024) of GALAH, and the DR3 catalogue of stellar parameters from Gaia RVS (Recio-Blanco et al., 2023). For APOGEE stars we require fe_h_flag = 0 and starflag=0, for GALAH we require flag_SP=0 and flag_fe_h=0, and for Gaia we require the first 13 flags of flags_gspspec to be zero. We also include in our comparison a compilation of measurements of metal-poor stars from the SAGA database (Suda et al., 2008)555Using version from April 7, 2021. Available at http://sagadatabase.jp/Data/txt_table_all.tsv..
@@
Figure 8 compares the metallicities measured by the two DESI pipelines with those of other surveys. The left (right) column refers to the RVS (SP) pipeline. Different rows show different surveys. We notice that there is a good correlation between DESI’s measurements and external measurements, however there are some clearly identifiable issues. The comparison with APOGEE clearly shows two sequences, one consistent with a small positive offset, and another with a negative offset, shared by the two pipelines. The SAGA sample shows a systematic bias in metallicities similarly shared by the two pipelines. The issue specific to the RVS pipeline seems to be a larger spread above the one-to-one line indicative of [Fe/H][Fe/H]external.
@@ 4.3 Radial velocity accuracy and precision
The quality of DESI radial velocity measurements has already been discussed in Cooper et al. (2023) and K24. The results of those papers should still apply to the survey validation part of the release. What we focus on here is the accuracy of radial velocities for the three programs of the main survey.
@@
Among the three programs of the main DESI survey, we highlight the fact that the backup survey is observed in conditions that are drastically worse than the bright survey (Dey et al., 2025). It has been discovered that these conditions (in particular, bright sky with the continuum solar-like spectrum) lead to large biases in the wavelength calibration code of the DESI pipeline (see Guy et al. (2023); DESI Collaboration et al. (2025)). These issues have been identified and rectified in the code of the DESI pipeline, but not in time for DR1. Because of that, the backup program data in DR1 show radial velocity systematics that are substantially worse than other programs, reaching levels of 20 km s-1 in about 1% of the footprint (see Appendix C for more details, and a simplistic position-dependent correction we suggest to use).
We now compare DESI radial velocities to those from Gaia RVS, APOGEE, and GALAH. For APOGEE and GALAH we apply the same quality cuts for the reference surveys as described in the previous section. For Gaia in addition to the previously mentioned cuts, we further require that the radial velocity error to be below 5 km s-1. On the DESI side, we apply the selection criteria listed at the beginning of Section 4 and require additionally that the radial velocity error is below 1 km s-1. Figure 11 shows the distribution of differences of DESI radial velocities with respect to external surveys or DESI’s own bright program. The bright program was not compared to GALAH due to the small number of overlapping stars – fewer than 50. Similarly, for the dark program, we compare only with DESI’s own measurements because the overlap with GALAH, APOGEE and Gaia RVS is negligible.
@@
The velocity offsets for the backup program are shown in the middle panel of Figure 11, and the distribution differs significantly from the bright program. Note the very significant tails of the distribution going all the way to 15 km s-1 caused by wavelength calibration issues of the backup program discussed earlier. While the median offsets for the backup program radial velocities are , , km s-1 for Gaia, APOGEE and GALAH, respectively, they are somewhat misleadingly low as the distribution is asymmetric with a long tail. A particularly large median offset with GALAH originates from the spatial structure of the velocity systematics in the backup survey (see Appendix C), caused by the fact that different locations on the sky were observed at different times and moon phases.
@@ 4.3.1 Radial velocity uncertainties
In Cooper et al. (2023), we have shown how the radial velocity error in DESI varies with colour and magnitude for MWS based on survey validation data. We have found, however, that for objects with typical colours , the radial velocity error mostly depends on the -band magnitude. The reason for this is that most of DESI’s radial velocity information comes from the Z arm of the instrument, due to its higher spectral resolution compared to bluer arms as well as higher total throughput in the red (DESI Collaboration et al., 2022). In Figure 12, we show the radial velocity uncertainty as computed by the RVS vs magnitude for stars in the bright program of the main survey. Here we use the same quality cuts as defined in the beginning of the section, limit the range of effective temperatures to K K (roughly corresponding to the colour range of ), and restrict the sample to those stars with effective exposure times close to the typical one of the bright survey (i.e. between 180 and 220 sec)666The effective exposure time for every source can be computed as TSNR2_LRG (see Guy et al., 2023, for more details), where the TSNR2_LRG column is available in the SCORES extension of the catalogue.. Since the radial velocity precision should be a function of metallicity, we split the plot between high metallicity dex (left panel) and low metallicity dex stars (right panel). As expected, high metallicity stars provide more precise radial velocities at the same magnitude due to a higher amount of spectral information caused by larger number and stronger absorption lines.
In the Figure we also overplot the approximate relation for the radial velocity error that incorporates the radial velocity precision change with metallicity. This relation was fitted to all the DESI stars in the bright program with :
We remark that the slope of the radial velocity error as a function of magnitude is slightly greater than 0.2. For background-dominated spectra, a slope of 0.4 is expected, while signal-dominated cases should yield a slope of 0.2. Thus the bright program is as expected closer to the signal-dominated case. The relation defined in Eq. 3 has been derived for the typical effective exposure time of the bright program, but there are sources in the catalogue with larger effective exposure times (for example if they have been observed multiple times over several passes). In this case the uncertainties should be scaled from Eq. 3 by where is effective exposure time for the source in seconds.
@@
We do not present a similar figure for the uncertainties of the [Fe/H] abundances since they are more affected by systematic errors. However, we find that the formal [Fe/H] error (for the RVS pipeline) can be similarly well described by a relation similar to Eqn 3:
@@ 4.3.2 Systematic floor of radial velocity measurements
Another important consideration for understanding the quality of DESI’s radial velocities is whether the reported radial velocity errors accurately reflect the true measurement uncertainties and whether any additional instrument systematic effects contribute to the radial velocity errors. One can approach this by analysing repeated radial velocity measurements of the same objects. Since in this VAC we include, for the first time, the measurements from single-epoch observations, we are now able to perform this analysis. Crucially, this type of test bypasses questions about absolute radial velocity offsets relative to ground truth (and whether such offsets vary from star to star), focusing instead solely on the repeatability of radial velocities and their associated uncertainties.
@@
A common approach to characterizing the repeatability is to assume that the calibrated radial velocity errors can be expressed as a function of the formal error through the following expression:
@@
where the factor scales the error and refers to a systematic floor (see, e.g., Simon & Geha, 2007; Li et al., 2019). If the data reduction pipeline correctly propagates all the uncertainties and the radial velocity measurement correctly takes them into account, then one expects . The systematic floor can be interpreted as an irreducible Gaussian noise added to every radial velocity measurement. The most common causes for that floor on various instruments are inaccuracies in the wavelength calibration or flexures of the spectrograph.
@@
It is possible to determine the constants and if we consider all pairs of observations for each star with radial velocities and uncertainties . Assuming a Gaussian distribution, and calibration of errors defined in Eqn. 5, one can then compute that should be normally distributed following .
@@
Motivated by this, in Figure 13 we show the robust estimate of the standard deviation of pair-wise radial velocity differences in bins of (see i.e. Koposov et al., 2011, for a similar approach). In the case of perfect errors and no systematic floor the points should lie along the y=x line.
@@
To ensure robustness against outliers (e.g. such as those introduced by binaries), the standard deviation of pairwise differences is estimated as half of the difference between 84-th and 16-th percentiles of the distribution. Grey points show measurements using all pairs of observations within the corresponding bin of , while black points include only pairs of observations separated by more than a day. The top, middle, and bottom panels show the results from bright, backup, and dark programs, respectively. As expected, the measured standard deviations flatten at small uncertainties (left side of the panels), indicating the presence of a systematic floor in radial velocity measurements. Additionally, the black points tend to lie above the grey points, as they include only pairs of observations separated by more than a day. This is expected if certain systematic errors are specific to a given night – for example, due to the particular arc calibrations used on that night. To determine the parameter and for different programs, we fit Eqn. 5 to the black data points shown in the Figure777The reason to fit for the systematic floor from black points that are above grey points is that we think the measurements of velocity differences from different nights provide a more honest assessment of systematics. In all cases, we have found that the best-fit value of the parameter (responsible for scaling of the uncertainties) is within a few percent of 1, hence we adopt . The values of the systematic floor measured for different programs are given in each panel of the Figure. We also show the best-fit models for each program on the Figure, demonstrating that the Eqn. 5 can well describe the observed behaviour.
@@
The systematic floor for the bright program is km (note that if we consider the floor measurement from all observations including observations on the same night the floor is km s-1, significantly better than 1 km -1). The floor for the backup program is around 2 km s-1; however, this value is somewhat misleading. As shown in Figure 11 and Appendix C, the backup program exhibits non-Gaussian systematic errors, making the formalism defined in Eqn. 5 inappropriate for characterizing its uncertainties. For the dark program the floor is found to be km s-1 which is higher than a systematic floor of km s-1 for the bright program program. The reason for this is not yet understood.
@@
As a last check of the calibrated radial velocity errors, we look at the distribution of pair-wise velocity differences divided by the combined calibrated error . We consider only observations separated by more than one day. If the calibrated errors are accurate, the distribution should be close to normal with a mean of zero and unit variance. This is shown in Figure 14 for three programs of the main survey. For both bright and dark programs, the distributions are close to a Gaussian, confirming that our calibration procedure is accurate. For the backup program, however, the distribution is clearly non-Gaussian, but this is expected due to the systematic issues discussed earlier.
@@
Finally, we note that the radial velocity errors reported in the VAC do not included the correction for the systematic floor, and users of the catalogue should apply it if it is appropriate for their science case.

### v3

@@ 4.1 Distribution of stellar parameters and abundances
Figure 5 shows the distribution of surface gravity vs (Kiel diagram) using measurements from both stellar pipelines. We show the stars observed in the main survey with signal-to-noise in the R arm of the instrument above 10. The main improvement in this distribution compared to the EDR catalogue (c.f. Figure 9 in K24) is the disappearance of clustering near the grid-points in and for the RVS catalogue. Otherwise, the distribution shows the expected main sequence, red giant branch as well as horizontal branch. For dwarf stars () with lower effective temperatures ( K), the surface gravity is not a smooth function of temperature, similar to what was seen in EDR. The morphology of the diagrams for RVS and SP pipeline is similar, but the SP parameters show some gridding of temperatures for cool stars; also, the RVS red giant branch seems less populated at .
@@
Finally, Figure 7 shows the metallicity distribution for stars in the main survey. We use the same selection of stars as for the previous figure. The distributions clearly show the peaks associated with the MW disk, and Gaia Sausage-Enceladus at [Fe/H] (see, e.g., Belokurov et al., 2018; Helmi et al., 2018; Feuillet et al., 2020). We note that although the metallicity distribution shows an almost power-law-like behaviour from [Fe/H] to lower metallicities, it also shows small "wiggles", for example at [Fe/H], where the metallicity distribution (as measured by two pipelines) briefly flattens. That could be the result of debris deposition from a relatively massive progenitor with the mean metallicity of (i.e. something like Thamnos or I’itoi; see Naidu et al., 2020). Overall the distribution of RVS and SP metallicities show similar behaviour, except near the edges. At high metallicities the RVS sample has more stars than SP (this is mainly caused by the exclusion of SP measurements fitted by PHOENIX models). At low metallicities we can see some differences in the SP and RVS distributions of [Fe/H] mostly due to the fact that the PHOENIX grid used by the RVS pipeline has a limit at [Fe/H], while the SP pipeline has a limit at [Fe/H]. We note that the SP metallicities show an excess near the grid edge at [Fe/H]= which is not physical. Using the same sample as shown in the figure, we can estimate the total number of extremely metal-poor stars (i.e with [Fe/H]) in the bright program of the main survey – it is if we use RVS abundances, and 3000 with SP abundances. These numbers rise by 2,000 if we include all surveys and programs.
@@ 4.2 Comparison of abundances to external measurements
In the previous section we looked at the overall distribution of stellar parameters and abundances. In this section we focus on iron abundance and compare our results to those from other large medium- and high-resolution surveys such as APOGEE (Majewski et al., 2017), GALAH (De Silva et al., 2015) and Gaia RVS (Cropper et al., 2018). A similar comparison to and Ca and Mg abundances is presented in the Appendix C. We use the data from the DR17 catalogue (Abdurro’uf et al., 2022) of APOGEE, the DR4 catalogue (Buder et al., 2024) of GALAH, and the DR3 catalogue of stellar parameters from Gaia RVS (Recio-Blanco et al., 2023). For APOGEE stars we require fe_h_flag = 0 and starflag=0, for GALAH we require flag_SP=0 and flag_fe_h=0, and for Gaia we require the first 13 flags of flags_gspspec to be zero. We also include in our comparison a compilation of measurements of metal-poor stars from the SAGA database (Suda et al., 2008).666Using version from April 7, 2021. Available at http://sagadatabase.jp/Data/txt_table_all.tsv.
@@
Figure 8 compares the metallicities measured by the two DESI pipelines with those of other surveys. The left (right) columns refer to the RVS (SP) pipeline. Different rows show different surveys. We notice that there is a good correlation between DESI’s measurements and external measurements, however there are some clearly identifiable issues. The comparison with APOGEE clearly shows two sequences, one consistent with a small positive offset, and another with a negative offset, shared by the two pipelines. The SAGA sample shows a systematic bias in metallicities similarly shared by the two pipelines. The issue specific to the RVS pipeline seems to be a larger spread above the one-to-one line indicative of [Fe/H][Fe/H]external.
@@ 4.3 Radial velocity accuracy and precision
The quality of DESI radial velocity measurements has already been discussed in Cooper et al. (2023) and K24. The results of those papers still apply to the survey validation part of the release. What we focus on here is the accuracy of radial velocities for the three programs of the main survey.
@@
Among the three programs of the main DESI survey, we highlight the fact that the backup survey is observed in conditions that are drastically worse than the bright survey (Dey et al., 2025). It has been discovered that these conditions (in particular, bright sky with the continuum solar-like spectrum) lead to large biases in the wavelength calibration code of the DESI pipeline (see Guy et al. (2023); DESI Collaboration et al. (2025)). These issues have been identified and rectified in the code of the DESI pipeline, but not in time for DR1. Because of that, the backup program data in DR1 show radial velocity systematics that are substantially worse than other programs, reaching levels of 20 km s-1 in about 1% of the footprint (see Appendix D for more details, and a simplistic position-dependent correction we suggest to use).
We now compare DESI radial velocities to those from Gaia RVS, APOGEE, and GALAH. For APOGEE and GALAH we apply the same quality cuts for the reference surveys as described in the previous section. For Gaia in addition to the previously mentioned cuts, we further require the radial velocity uncertainty to be below 5 km s-1. On the DESI side, we apply the selection criteria listed at the beginning of Section 4 and require additionally that the radial velocity uncertainty is below 1 km s-1. Figure 11 shows the distribution of differences of DESI radial velocities with respect to external surveys or DESI’s own bright program. The bright program was not compared to GALAH due to the small number of overlapping stars – fewer than 50. Similarly, for the dark program, we compare only with DESI’s own measurements because the overlap with GALAH, APOGEE and Gaia RVS is negligible.
@@
The velocity offsets for the backup program are shown in the middle panel of Figure 11, and the distribution differs significantly from the bright program. Note the very significant tails of the distribution going all the way to 15 km s-1 caused by wavelength calibration issues of the backup program discussed earlier. While the median offsets for the backup program radial velocities are , , km s-1 for Gaia, APOGEE and GALAH, respectively, they are somewhat misleadingly low as the distribution is asymmetric with a long tail. A particularly large median offset with GALAH originates from the spatial structure of the velocity systematics in the backup survey (see Appendix D), caused by the fact that different locations on the sky were observed at different times and moon phases.
@@ 4.3.1 Radial velocity uncertainties
In Cooper et al. (2023), we have shown how the radial velocity uncertainty in DESI varies with colour and magnitude for MWS based on survey validation data. We have found, however, that for objects with typical colours , the radial velocity uncertainty mostly depends on the -band magnitude. The reason for this is that most of DESI’s radial velocity information comes from the Z arm of the instrument, due to its higher spectral resolution compared to bluer arms as well as higher total throughput in the red (DESI Collaboration et al., 2022). In Figure 12, we show the radial velocity uncertainty as computed by the RVS vs magnitude for stars in the bright program of the main survey. Here we use the same quality cuts as defined in the beginning of the section, limit the range of effective temperatures to K K (roughly corresponding to the colour range of ), and restrict the sample to those stars with effective exposure times close to the typical one of the bright survey (i.e. between 180 and 220 sec).777The effective exposure time for every source can be computed as TSNR2_LRG (see Guy et al., 2023, for more details), where the TSNR2_LRG column is available in the SCORES extension of the catalogue. Since the radial velocity precision should be a function of metallicity, we split the plot between high metallicity dex (left panel) and low metallicity dex stars (right panel). As expected, high metallicity stars provide more precise radial velocities at the same magnitude due to a higher amount of spectral information caused by larger number and stronger absorption lines.
In the Figure we also overplot the approximate relation for the radial velocity uncertainty that incorporates the radial velocity precision change with metallicity. This relation was fitted to all the DESI stars in the bright program with :
We remark that the slope of the radial velocity uncertainty as a function of magnitude is slightly greater than 0.2. For background-dominated spectra, a slope of 0.4 is expected, while signal-dominated cases should yield a slope of 0.2. Thus the bright program is as expected closer to the signal-dominated case. The relation defined in Eq. 3 has been derived for the typical effective exposure time of the bright program, but there are sources in the catalogue with larger effective exposure times (for example if they have been observed multiple times over several passes). In this case the uncertainties need to be scaled from Eq. 3 by where is effective exposure time for the source in seconds.
@@
We do not present a similar figure for the uncertainties of the [Fe/H] abundances since they are more affected by systematic errors. However, we find that the formal [Fe/H] uncertainty (for the RVS pipeline) can be similarly well described by a relation similar to Eqn 3:
@@ 4.3.2 Systematic floor of radial velocity measurements
Another important consideration for understanding the quality of DESI’s radial velocities is whether the reported radial velocity uncertainties accurately reflect true measurement uncertainties and whether any additional instrument systematic effects contribute to the radial velocity errors. One can approach this by analysing repeated radial velocity measurements of the same objects. Since in this VAC we include, for the first time, the measurements from single-epoch observations, we are now able to perform this analysis. Crucially, this type of test bypasses questions about absolute radial velocity offsets relative to ground truth (and whether such offsets vary from star to star), focusing instead solely on the repeatability of radial velocities and their associated uncertainties.
@@
A common approach to characterizing the repeatability is to assume that the calibrated radial velocity uncertainties can be expressed as a function of the formal uncertainty through the following expression:
@@
where the factor scales the uncertainty and refers to a systematic floor (see, e.g., Simon and Geha, 2007; Li et al., 2019). If the data reduction pipeline correctly propagates all the uncertainties and the radial velocity measurement correctly takes them into account, then one expects . The systematic floor can be interpreted as an irreducible Gaussian noise added to every radial velocity measurement. The most common causes for that floor on various instruments are inaccuracies in the wavelength calibration or flexures of the spectrograph.
@@
It is possible to determine the constants and if we consider all pairs of observations for each star with radial velocities and uncertainties . Assuming a Gaussian distribution, and calibration of uncertainties defined in Eqn. 5, one can then compute that should be normally distributed following .
@@
Motivated by this, in Figure 13 we show the robust estimate of the standard deviation of pair-wise radial velocity differences in bins of (see i.e. Koposov et al., 2011, for a similar approach). In the case of perfect uncertainties and no systematic floor the points should lie along the y=x line.
@@
To ensure robustness against outliers (e.g. such as those introduced by binaries), the standard deviation of pairwise differences is estimated as half of the difference between 84-th and 16-th percentiles of the distribution. Grey points show measurements using all pairs of observations within the corresponding bin of , while black points include only pairs of observations separated by more than a day. The top, middle, and bottom panels show the results from bright, backup, and dark programs, respectively. As expected, the measured standard deviations flatten at small uncertainties (left side of the panels), indicating the presence of a systematic floor in radial velocity measurements. Additionally, the black points tend to lie above the grey points, as they include only pairs of observations separated by more than a day. This is expected if certain systematic errors are specific to a given night – for example, due to the particular arc calibrations used on that night. To determine the parameters and for different programs, we fit Eqn. 5 to the black data points shown in the Figure.888The reason to fit for the systematic floor from black points that are above grey points is that we think the measurements of velocity differences from different nights provide a more honest assessment of systematics In all cases, we have found that the best-fit value of the parameter (responsible for scaling of the uncertainties) is within a few percent of 1, hence we adopt . The values of the systematic floor measured for different programs are given in each panel of the Figure. We also show the best-fit models for each program on the Figure, demonstrating that the Eqn. 5 can well describe the observed behaviour.
@@
The systematic floor for the bright program is km s-1 (note that if we consider the floor measurement from all observations including observations on the same night the floor is km s-1, significantly better than 1 km -1). The floor for the backup program is around 2 km s-1; however, this value is somewhat misleading. As shown in Figure 11 and Appendix D, the backup program exhibits non-Gaussian systematic errors, making the formalism defined in Eqn. 5 inappropriate for characterizing its uncertainties. For the dark program the floor is found to be km s-1 which is higher than a systematic floor of km s-1 for the bright program. The reason for this is not yet understood.
@@
As a last check of the calibrated radial velocity uncertainties, we look at the distribution of pair-wise velocity differences divided by the combined calibrated uncertainty . We consider only observations separated by more than one day. If the calibrated uncertainties are accurate, the distribution should be close to normal with a mean of zero and unit variance. This is shown in Figure 14 for three programs of the main survey. For both bright and dark programs, the distributions are close to a Gaussian, confirming that our calibration procedure is accurate. For the backup program, however, the distribution is clearly non-Gaussian, but this is expected due to the systematic issues discussed earlier.
@@
Finally, we note that the radial velocity uncertainties reported in the VAC do not include the correction for the systematic floor, and users of the catalogue should apply it if it is appropriate for their science case.